# Parameterization-driven Neural Implicit Surfaces Editing

### v2

@@ Abstract
The growing capabilities of neural rendering have increased the demand for new techniques that enable intuitive editing of 3D objects, particularly when they are represented as neural implicit surfaces. In this paper, we present a novel neural algorithm to parameterize neural implicit surfaces to simple parametric domains, such as spheres, cubes, or polycubes, thereby facilitating visualization and various editing tasks. Technically, our method computes a bi-directional deformation between 3D objects and their chosen parametric domains, eliminating the need for any prior information. We adopt a forward mapping of points on the zero level set of the 3D object to a parametric domain, followed by a backward mapping through inverse deformation. To ensure the map is bijective, we employ a cycle loss while optimizing the smoothness of both deformations. Additionally, we leverage a Laplacian regularizer to effectively control angle distortion and offer the flexibility to choose from a range of parametric domains for managing area distortion. Designed for compatibility, our framework integrates seamlessly with existing neural rendering pipelines, taking multi-view images as input to reconstruct 3D geometry and compute the corresponding texture map. We also introduce a simple yet effective technique for intrinsic radiance decomposition, facilitating both view-independent material editing and view-dependent shading editing. Our method allows for the immediate rendering of edited textures through volume rendering, without the need for network re-training. Moreover, our approach supports the co-parameterization of multiple objects and enables texture transfer between them. We demonstrate the effectiveness of our method on images of human heads and man-made objects. We will make the source code publicly available.
Figure 1: Overview of NeRF-based editing work. We summarize the existing work into three main categories: manipulating the latent codes, re-training, and parameterization. In this study, we propose a novel NeRF editing method based on our 3D neural parameterization framework.
@@ 1 Introduction
Neural radiance fields (NeRF) [29] have garnered remarkable success in both the computer vision and computer graphics communities, redefining benchmarks for high-quality renderings and novel view synthesis. Building upon NeRF, a variety of methods leveraging implicit neural representations [32, 47, 61] have emerged, delivering high-fidelity 3D reconstructions. The demand for intuitive 3D object editing techniques has increased with the growing capabilities of neural rendering, especially for neural implicit surfaces. Extensive explorations have been made in editing within the Neural Rendering framework to achieve diverse visual effects. By manipulating the latent codes of the network, one can implicitly alter the shape and color of 3D objects[25, 63, 54]. In the work of [45, 20], palette-based methods have been employed to facilitate transformations in the color and style of rendered scenes. However, these methods suffer from limited controllability and disallow local pixel editing. Seal3D [48] proposed a teacher-student training strategy to edit the interactive scene at the pixel level. Yet, this method is tricky to handle shading, and each subsequent edit required a new training iteration. By employing neural networks to formulate a differentiable surface parameterization, it can be facilitated to edit textures or shapes within the neural rendering pipeline [53, 27]. Existing neural parameterization techniques still face significant challenges, such as the need for appropriate distortion constraints to ensure accurate detail reconstruction or reducing the reliance on prior information from tracked mesh and UV mapping. The final quality of edited visual effects is significantly impacted by these issues. We summarize the main work related to NeRF-editing in Fig. 1.
@@
This paper aims at developing an intuitive and easy-to-use tool for appearance editing in neural implicit surfaces. Towards this goal, we propose a novel method to parameterize neural implicit surfaces to parametric domains where the radiance field is represented as a texture map. Distinguished from previous work that using 2D texture mapping, our neural parameterization algorithm supports the selection of different 3D parametric domains according to geometric shapes to avoid seams and reduce distortion, such as spheres, cubes, and polycubes. Technically, our parameterization algorithm utilizes a neural network to learn a bi-directional deformation between 3D objects and the chosen parametric domains. This involves a forward deformation that maps points from the zero level set of the neural implicit surface to the parametric domain, followed by an inverse deformation, mapping points backward. Notably, we do not require any explicit prior information in learning both deformations. We employ a cycle loss to ensure the smoothness of the bi-directional deformation, and a Laplacian regularization to effectively control angle distortion. Our method also supports the flexible selection of parametric domains for controlling area distortion. With the parametrization, 3D radiance field is mapped to a simple 3D parametric domain which can be further disassembled into a 2D domain, facilitating visualization and various editing tasks. Furthermore, we decompose the 2D radiance field into two components: view-independent material and view-dependent shading, streamlining both texture and shading editing. Our neural parameterization algorithm is fully compatible with existing neural rendering pipelines, allowing 3D reconstruction from multi-view images as input and the creation of texture maps simultaneously. It allows for the immediate rendering of edited textures through volume rendering, without network re-training. Moreover, it supports co-parameterization of objects of similar geometry and enables texture transfer between them. We validate the effectiveness of our method on scenes of human heads and man-made objects. Our contributions are summarized as follows:
@@
  • We present a neural parameterization framework that computes bi-directional deformation between 3D objects and their parametric domains, eliminating the need for any prior information from tracked mesh or UV mapping. Our approach utilizes Laplacian loss to minimize angle distortion and provides a choice of 3D parametric domains, such as spheres, cubes, or polycubes, to control area distortion.
@@
  • Our neural parameterization framework is fully compatible with existing neural rendering architectures, accepting multi-view images as input to reconstruct 3D geometry and generate UV maps as output. Additionally, it enables the direct rendering of modified textures using the volume rendering pipeline.
@@ 2 Related Work
Neural Implicit Functions. Recent years have seen a surge in the successful application of neural implicit functions within the realm of 3D deep learning. Compared to explicit representations such as point clouds [37], voxels [36], and meshes [46], neural implicit representations offer unique advantages including flexibility, continuity, and robustness [28, 40, 33]. When integrated within the neural rendering pipeline, neural implicit surfaces exhibit exceptional quality in 3D reconstructions from multi-view images [21, 49, 50, 54, 30, 22, 38]. However, as geometries and radiance fields are encoded as network parameters, editing them becomes complex and non-intuitive, compared to explicit representations. This paper aims at tackling this challenge by explicitly representing radiance fields as texture mapping that encompasses both view-dependent and view-independent components.
NeRF Editing. Numerous studies have focused on editing neural radiance fields, including relighting [42, 23], composition [34, 56], content generation [31], and shape editing [65, 24]. Editing tools operating at the scene- or object-level, such as [56, 63, 16], utilize latent codes to modify and stylize the appearance of an object or entire scenes. Conversely, pixel-level editing tools, such as NeuMesh [57] and Seal3D [48], utilize training-based approaches to edit fine-grained details. However, NeuMesh relies on explicit meshes which lack efficiency and robustness. Although Seal3D achieves interactive editing, it still suffers from poor geometry reconstruction and cannot change the illumination scene. Parameterization-based methods, such as NeuTex [53], ISO-UVField [39] and NeP [27], integrate the learning of UV mapping into the neural rendering framework, providing an intuitive means for editing within 2D domains. However, these methods often suffer from large distortions in the parametrization. Furthermore, they require 3D prior information from tracked meshes and UV mapping - conditions that may not be readily met in practical, real-world scenarios. IntrinsicNeRF [62] enhances standard neural radiance fields by generating additional outputs, including reflectance, shading, and a residual term. Incorporating a semantic branch, it facilitates real-time scene editing. However, it often results in aggregated colors, akin to palette-based methods [45, 20], hence cannot offer fine-grained editing results at the pixel level. Our approach decomposes the radiance field into view-dependent shadings and view-independent materials. Both components are represented as textures, enabling intuitive material modifications while preserving consistent shading.
@@ 3 Method
The input of our method is a collection of RGB images, denoted =$\mathcal{I}=\{\mathbf{I}_{i}\}$, representing either multiple objects of similar geometry or a single object, captured from different viewpoints. To differentiate among objects, we assign a latent shape code $\mathbf{z}_{s}$ and an appearance code $\mathbf{z}_{a}$ to each object. This facilitates the co-parameterization of multiple objects and enables the transfer of material and shading among them. As illustrated in Figure 2, our method reconstructs 3D geometry as a neural implicit surface and encodes the 3D radiance field in a simple parametric domain, such as spheres and polycubes. For each input image, we randomly select a set of pixels. Then, for each pixel, we shoot a ray originating from the location of camera and traversing through the pixel. If the ray intersects with the surface, at a point, denoted as $\mathbf{p}$, we compute its corresponding position in the parametric domain $\mathcal{D}$, denoted as $\mathbf{p}^{\prime}$, via a bi-directional deformation (Sec. 3.1): the forward deformation (depicted in orange) generates a displacement vector so that the updated position $\mathbf{p}^{\prime}$ is in the parametric domain; subsequently, the inverse deformation (depicted in green) maps $\mathbf{p}^{\prime}$ back to a point $\mathbf{p}^{\prime\prime}$, which is anticipated to be in close proximity to the surface. We incorporate a cycle loss to penalize occurrences where $\mathbf{p}^{\prime\prime}$ deviates from $\mathbf{p}$. In addition, we employ a Laplacian loss (Sec. 3.2) to effectively reduce the angle distortion between the parametric domain and the original surface. Furthermore, we decompose the radiance into a view-independent material field and a view-dependent shading field (Sec. 3.3), each of which can be edited independently, thus augmenting the editing capability of our framework.
GT Normal $\mathcal{D}$ $\mathcal{S}$
Figure 3: Parameterization results. Human heads are co-parameterized to a sphere owing to the clear geometric resemblance (top four rows). Given the diverse geometry of cars, each is parameterized to a polycube domain (bottom three rows). From left to right: the input image, the normal map of the reconstructed geometry $\mathcal{S}$, the parametric domain $\mathcal{D}$, and the reconstructed surface $\mathcal{S}$. See supplementary material for additional results.
@@ 3.1 Bi-directional deformation
The bi-directional deformation module serves as the pivotal component within our neural parameterization framework. As illustrated in Figure 2, this module consists of two sub-networks: $F_{\text{def}}$, which is responsible for the forward mapping from the original model to the parameter domain, and $F_{\text{inv-def}}$, which establishes the inverse mapping.
@@
Remark. It is worth mentioning that while our framework involves two SDFs - one for the 3D surface $\mathcal{S}$ and the other for the parametric domain $\mathcal{D}$ our network design only requires one of them. This is because one SDF can be derived from the other via either the forward or the backward deformation. To reduce the network complexity, we only adopt one geometry sub-network $F_{\text{sdf}}$, which is used to represent the parametric domain $\mathcal{D}$. Given a sphere, a cube, or a polycube, we pretrain the SDF sub-network to encode $\mathcal{D}$’s geometry. Leveraging the inherent smoothness of MLPs, the resulted SDF is smooth, aiding the computation of Laplacian.
@@ 3.3 Appearance Decomposition
Radiance represents the directional emission of color, illustrating how the appearance of an object varies in response to changes in the observational viewpoint [29]. Drawing inspiration from previous works [27, 62] and applying the assumptions of Lambertian and grayscale shading [4], we decompose the radiance field of the target surface into two components: a view-independent material field and a view-dependent shading field (see Figure 9 (b)). This decomposition allows for independent editing of material and shading. More precisely, for a sample point $\mathbf{p}\in\mathcal{S}$, we employ two sub-networks, $F_{\text{mat}}$ and $F_{\text{shd}}$, to decompose the radiance $\mathbf{r}\in\mathbb{R}^{3}$ at point $\mathbf{p}$ as follows:
@@
Then we employ the SDF-based volume rendering [61] to integrate the radiances along the ray to compute color $\mathbf{c}_{\text{pred}}=\sum_{i}\lambda_{i}\mathbf{r}_{i}$, where $\mathbf{r}_{i}$ is the radiance of a sample point $\mathbf{p}_{i}$ and $\lambda_{i}$ is the same color weight as used in Equation (4).
Refer to captionRefer to captionRefer to captionRefer to caption (a) $\mathcal{S}$ (b) 2.505 (c) 6.094 (d) 1.633Refer to captionRefer to captionRefer to captionRefer to caption (a) $\mathcal{S}$ (b) 1.793 (c) 1.932 (d) 1.425
Figure 4: We offer the flexibility to choose from a range of parametric domains for managing area distortion. The closer the parametric domain $\mathcal{D}$((b-d)) to the (a) 3D surface $\mathcal{S}$, the lower the area distortion $\mathcal{E}_{\text{area}}$ (shown below the objects, with a detailed definition provided in Section 4) in the parameterization.
3.4 Choice of Parametric Domains
An ideal parameterization should be conformal and area-preserving to avoid any signal distortion. For general objects, the overall parameterization to 2D domains, such as disks or squares, leads to large distortion. Thus, the standard 2D texture mapping is to cut the surface into several patches, each of which can be parameterized with low distortion. However, this multi-chart approach inevitably produces seams, and the process of creating a good overall mapping also requires one to define suitable boundaries of the charts in the 2D domain. Therefore, unlike NeP [27], we map the neural implicit surfaces to 3D parametric domains and provide the selection of an appropriate domain shape to reduce area distortion. In addition, we use the Laplacian regularizer to reduce angle distortion. Given the existence of infinite conformal maps [18], identifying those with minimal area distortion is highly desired. Area distortion is influenced by the similarity between the 3D surface and the parametric domain: the closer their alignment, the less the distortion. Thus, a simplistic, closely aligned parametric domain is ideal, balancing simplicity and alignment. Our algorithm accommodates various domains such as spheres, cubes, and polycubes, allowing flexibility in selection. It should be noted that our method supports more complex parametric domains, such as polyhedral complexes [59], thanks to the representation capability of the SDF sub-network. In our current implementation, we use spheres, cubes and polycubes, mainly because of their simplicity. Figure 4 illustrates the impact on the choice of parametric domain to the parameterization quality.

### v3

# Parameterization-driven Neural Surface Reconstruction for Object-oriented Editing in Neural Rendering

@@ Abstract
The advancements in neural rendering have increased the need for techniques that enable intuitive editing of 3D objects represented as neural implicit surfaces. This paper introduces a novel neural algorithm for parameterizing neural implicit surfaces to simple parametric domains like spheres and polycubes. Our method allows users to specify the number of cubes in the parametric domain, learning a configuration that closely resembles the target 3D object’s geometry. It computes bi-directional deformation between the object and the domain using a forward mapping from the object’s zero level set and an inverse deformation for backward mapping. We ensure nearly bijective mapping with a cycle loss and optimize deformation smoothness. The parameterization quality, assessed by angle and area distortions, is guaranteed using a Laplacian regularizer and an optimized learned parametric domain. Our framework integrates with existing neural rendering pipelines, using multi-view images of a single object or multiple objects of similar geometries to reconstruct 3D geometry and compute texture maps automatically, eliminating the need for any prior information. We demonstrate the method’s effectiveness on images of human heads and man-made objects. The source code is available at https://xubaixinxbx.github.io/neuparam.
Keywords:
Polycube parameterization Neural surface reconstruction and editing Neural rendering
footnotetext: Corresponding author: Ying He (email: yhe@ntu.edu.sg).
@@ 1 Introduction
Neural radiance fields (NeRF) [29] have garnered remarkable success in both the computer vision and computer graphics communities, redefining benchmarks for high-quality renderings and novel view synthesis. Building upon NeRF, a variety of methods leveraging implicit neural representations [32, 49, 61] have emerged, delivering high-fidelity 3D reconstructions. The demand for intuitive 3D object editing techniques has increased with the growing capabilities of neural rendering, especially for neural implicit surfaces. Extensive explorations have been made in editing within the Neural Rendering framework to achieve diverse visual effects. By manipulating the latent codes of the network, one can implicitly alter the shape and color of 3D objects[25, 64, 56]. In the work of [47, 20], palette-based methods have been employed to facilitate transformations in the color and style of rendered scenes. However, these methods suffer from limited controllability and disallow local pixel editing. Seal3D [50] proposed a teacher-student training strategy to edit the interactive scene at the pixel level. Yet, this method is tricky to handle shading, and each subsequent edit required a new training iteration. By employing neural networks to formulate a differentiable surface parameterization, it can be facilitated to edit textures or shapes within the neural rendering pipeline [55, 27]. Existing neural parameterization techniques still face significant challenges, such as the need for appropriate distortion constraints to ensure accurate detail reconstruction or reducing the reliance on prior information from tracked mesh and UV mapping. The final quality of edited visual effects is significantly impacted by these issues.
@@
This paper aims to develop an intuitive and easy-to-use tool for appearance editing in neural implicit surfaces within the neural rendering framework, utilizing only 2D images as input. Towards this goal, we propose a novel method to parameterize neural implicit surfaces to parametric domains where the radiance field is represented as a texture map. Distinguished from previous work that using 2D texture mapping, our neural parameterization algorithm supports the selection of different 3D parametric domains according to geometric shapes to avoid seams and reduce distortion, such as spheres, and polycubes. Technically, our parameterization algorithm utilizes learnable parameters to fit a parametric domain first and a neural network to learn a bi-directional deformation between 3D objects and the chosen parametric domains. This involves a forward deformation that maps points from the zero level set of the neural implicit surface to the parametric domain, followed by an inverse deformation, mapping points backward. Notably, we do not require any explicit prior information in learning both deformations. We employ a cycle loss to ensure the smoothness of the bi-directional deformation, and a Laplacian regularization to effectively control angle distortion. With the parameterization, 3D radiance field is mapped to a simple 3D parametric domain which can be further disassembled into a 2D domain, facilitating visualization and various editing tasks. Furthermore, we decompose the 2D radiance field into two components: view-independent material and view-dependent shading, streamlining both texture and shading editing. Our neural parameterization algorithm is fully compatible with existing neural rendering pipelines, allowing 3D reconstruction from multi-view images as input and the creation of texture maps simultaneously. It allows for the immediate rendering of edited textures through volume rendering, without network re-training. Moreover, it supports co-parameterization of objects of similar geometry and enables texture transfer between them. We validate the effectiveness of our method on scenes of human heads and man-made objects. Our contributions are summarized as follows:
@@
  • We present a neural parameterization framework that automatically learns a parametric domain and computes bi-directional deformation between 3D objects and their parametric domains, eliminating the need for any prior information from tracked mesh or UV mapping. Our approach utilizes Laplacian loss to minimize angle distortion and allows the user to specify the domain complexity to control area distortion.
@@
  • Our neural parameterization framework is end-to-end and fully compatible with existing neural rendering architectures, accepting multi-view images as input to reconstruct 3D geometry and generate UV maps as output. Additionally, it enables the direct rendering of modified textures using the volume rendering pipeline.
@@ 2 Related Work
Neural Implicit Functions. Recent years have seen a surge in the successful application of neural implicit functions in 3D deep learning. Compared to explicit representations such as point clouds [36], voxels [37], and meshes [48], neural implicit representations offer unique advantages including flexibility, continuity, and robustness [28, 41, 33]. When integrated within the neural rendering pipeline, neural implicit surfaces exhibit exceptional quality in 3D reconstructions from multi-view images [21, 51, 52, 56, 30, 22, 39]. However, as geometries and radiance fields are encoded as network parameters, editing them becomes complex and non-intuitive, compared to explicit representations. This paper aims at tackling this challenge by explicitly representing radiance fields as texture mapping that encompasses both view-dependent and view-independent components.
NeRF Editing. Numerous studies have focused on editing neural radiance fields, including relighting [44, 23], composition [34, 58], content generation [31], and shape editing [65, 24]. Editing tools operating at the scene- or object-level, such as [58, 64, 17], utilize latent codes to stylize the appearance of an object or entire scenes. Conversely, pixel-level editing tools, such as NeuMesh [57] and Seal3D [50], utilize training-based approaches to edit fine-grained details. However, NeuMesh relies on explicit meshes which lack efficiency and robustness. Although Seal3D achieves interactive editing, it still suffers from poor geometry reconstruction and cannot change the illumination scene. Parameterization-based methods, such as NeuTex [55], ISO-UVField [40] and NeP [27], integrate the learning of UV mapping into neural rendering, enabling intuitive editing within 2D domains. However, these methods often suffer from large distortions in the parametrization. Furthermore, they require 3D prior information from tracked meshes and UV mapping - conditions that may not be readily met in practical, real-world scenarios. IntrinsicNeRF [63] enhances standard neural radiance fields by generating additional outputs, including reflectance, shading, and a residual term. Incorporating a semantic branch, it facilitates real-time scene editing. However, it often results in aggregated colors, akin to palette-based methods [47, 20], hence cannot offer fine-grained editing results at the pixel level. We decompose the radiance field into view-dependent shadings and view-independent materials. Both components are represented as textures, enabling intuitive material modifications while preserving consistent shading.
@@ 3 Method
The input of our method is a collection of RGB images, denoted =$\mathcal{I}=\{\mathbf{I}_{i}\}$, representing multiple objects of similar geometry or a single object, captured from different viewpoints. To differentiate among objects, we assign a latent shape code $\mathbf{z}_{s}$ and an appearance code $\mathbf{z}_{a}$ to each object. This facilitates the co-parameterization of multiple objects and enables the transfer of material and shading among them. As illustrated in Fig. 1, our method reconstructs 3D geometry as a neural implicit surface and encodes the 3D radiance field in a simple parametric domain, such as spheres and polycubes, which we learn from an initial phase in our training (Sec. 3.1). For each input image, we randomly select a set of pixels. Then, for each pixel, we shoot a ray originating from the location of camera and traversing through the pixel. If the ray intersects with the surface, at a point, denoted as $\mathbf{p}$, we compute its corresponding position in the parametric domain $\mathcal{D}$, denoted as $\mathbf{p}^{\prime}$, via a bi-directional deformation (Sec. 3.2): the forward deformation (depicted in orange) generates a displacement vector so that the updated position $\mathbf{p}^{\prime}$ is in the parametric domain; subsequently, the inverse deformation (depicted in green) maps $\mathbf{p}^{\prime}$ back to a point $\mathbf{p}^{\prime\prime}$, which is anticipated to be in close proximity to the surface. We incorporate a cycle loss to penalize occurrences where $\mathbf{p}^{\prime\prime}$ deviates from $\mathbf{p}$. In addition, we employ a Laplacian loss (Sec. 3.3) to effectively reduce the angle distortion between the parametric domain and the original surface. Furthermore, we decompose the radiance into a view-independent material field and a view-dependent shading field (Sec. 3.4), each of which can be edited independently, thus augmenting the editing capability of our framework. See Fig. 1 for the algorithmic pipeline.
3.1 Parametric Domain Learning
In our framework, we consider two types of parametric domains: spheres and polycubes, both represented as signed distance fields. Spheres are characterized by four learnable parameters: the center position and radius. Polycubes, a generalization of cubes, are simple structures yet offer sufficient degrees of freedom to capture the salient geometric features of target surfaces. We define a learnable polycube with $k$ elements as follows: For the $i$-th element, which is an axis-aligned box, its SDF $\phi_{i}$ is given by
$\phi_{i}=\max(|x-x_{i}|-a_{i},|y-y_{i}|-b_{i},|z-z_{i}|-c_{i}),$ (1)
where $(x_{i},y_{i},z_{i})$ represents the box center, and $(a_{i},b_{i},c_{i})$ is its dimensions. By combining the $k$ boxes, the SDF of the polycube is obtained as
$f_{\text{pc}}=\max(-\phi_{1},-\phi_{2},...,-\phi_{k}).$ (2)
To ensure the differentiability of Eq. (2), we approximate it using Kreisselmeier-Steinhauser (KS) function [19] as
$f_{\text{pc}}\approx-1/\lambda*\log\left(\sum_{i}\exp(-\lambda*\phi_{i})\right),$
where $\lambda$ is a typically large value, set to 100 in our experiments.
Since determining the configuration of boxes does not require a highly accurate geometry of the target object $\mathcal{S}$, we design the parametric domain learning as a one-time, intermediate step in the neural rendering pipeline. Specifically, after an initial training phase of approximately 100 epochs, the coarse geometry of the surface $\mathcal{S}$ is obtained. Subsequently, we activate the domain learning module, which takes the SDF $G_{\text{sdf}}$ of the coarse surface $\mathcal{S}$ and the user-specified number of boxes $k$ as input. It computes the SDF $F_{\text{sdf}}$ of the parametric domain $\mathcal{D}$ via minimizing the loss function:
$\mathcal{L}_{\text{param}}=\mathcal{L}_{\text{sdf}}+\lambda_{s}\mathcal{L}_{\text{lapsdf}},$ (3)
where
$\mathcal{L}_{\text{sdf}}=\left|F_{\text{sdf}}(p)-G_{\text{sdf}}(p)\right|,\mathcal{L}_{\text{lapsdf}}=\left|\frac{\partial^{2}F_{\text{sdf}}}{\partial x^{2}}+\frac{\partial^{2}F_{\text{sdf}}}{\partial y^{2}}+\frac{\partial^{2}F_{\text{sdf}}}{\partial z^{2}}\right|,$ (4)
and the weight $\lambda_{s}$ is set to 0.01. The term $\mathcal{L}_{\text{lapsdf}}$, which is the Laplacian of the signed distance field of the polycube $\mathcal{D}$, plays a critical role in controlling its geometric complexity.
$\mathcal{S}$ Iter 0 Iter 5
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Iter 10 Iter 45 Iter 55
After obtaining the SDF $F_{\text{sdf}}$ of the polycube $\mathcal{D}$, we deprecate the SDF network $G_{\text{sdf}}$ and instead use the deformation network $F_{\text{def}}$ and parametric domain SDF $F_{\text{sdf}}$ network to represent the source surface $\mathcal{S}$. The right inset showcases a polycube learning example for a human head model. Given a user-specified parameter $k=3$, our method initializes three randomly placed boxes. It proceeds to minimize the loss function $\mathcal{L}_{\text{param}}$, iteratively refining the size and positioning of each box to ensure the polycube closely approximates the surface $\mathcal{S}$. This parametric domain learning phase typically converges within 30 to 60 iterations, taking under 10 minutes to complete. For more details, see the supplementary video.
@@ 3.2 Bi-directional Deformation
The bi-directional deformation module serves as the pivotal component within our neural parameterization framework. As illustrated in Fig. 1, this module consists of two sub-networks: $F_{\text{def}}$, which is responsible for the forward mapping from the original model to the parameter domain, and $F_{\text{inv-def}}$, which establishes the inverse mapping.
@@ 3.4 Appearance Decomposition
Radiance represents the directional emission of color, illustrating how the appearance of an object varies in response to changes in the observational viewpoint [29]. Drawing inspiration from previous works [27, 63] and applying the assumptions of Lambertian and grayscale shading [5], we decompose the radiance field on the parametric domain into two components: a view-independent material field and a view-dependent shading field (see Fig. 6(a)). This decomposition allows for independent editing of material and shading. More precisely, for a sample point $\mathbf{p}\in\mathcal{S}$, we employ two sub-networks, $F_{\text{mat}}$ and $F_{\text{shd}}$, to decompose the radiance $\mathbf{r}\in\mathbb{R}^{3}$ at point $\mathbf{p}$ as follows:
@@
Then we employ the SDF-based volume rendering [61] to integrate the radiances along the ray to compute color $\mathbf{c}_{\text{pred}}=\sum_{i}\lambda_{i}\mathbf{r}_{i}$, where $\mathbf{r}_{i}$ is the radiance of a sample point $\mathbf{p}_{i}$ and $\lambda_{i}$ is the same color weight as used in Eq. (8).